cotxdy=ydx=(cotx)(3e^sinx)dx


1
Expert's answer
2021-05-17T14:56:14-0400

If cotxdy=ydx=(cotx)(3e^sinx)dx then "\left\{ \begin{array}{l}
ydx = \cot xdy\\
ydx = \cot x \cdot 3{e^{\sin x}}dx
\end{array} \right."

From 1-st equation:

"\frac{{dy}}{y} = \frac{{\sin x}}{{\cos x}}dx = - \frac{{d\cos x}}{{\cos x}} \Rightarrow \ln y = - \ln \cos x + \ln {C_1} = \ln \frac{{{C_1}}}{{\cos x}} \Rightarrow y = \frac{{{C_1}}}{{\cos x}}"

From 2-nd equation:


"ydx = \cot x \cdot 3{e^{\sin x}}dx \Rightarrow \frac{{{C_1}}}{{\cos x}}dx = \cot x \cdot 3{e^{\sin x}}dx \Rightarrow \frac{{{C_1}\sin x}}{{\cos x}}dx = \cos x \cdot 3{e^{\sin x}}dx \Rightarrow - {C_1}\frac{{d\cos x}}{{\cos x}} = 3{e^{\sin x}}d\sin x \Rightarrow - {C_1}\ln \cos x = 3{e^{\sin x}} + {C_2} \Rightarrow {C_1}\ln \frac{1}{{\cos x}} = 3{e^{\sin x}} + {C_2}"

The resulting equality will not hold for any "{C_1}, {C_2}" .

So, system has no solutions.

Answer: There are no solutions.